OpenStudy (anonymous):

5C + 2SO----->CS2+4CO A. How many moles of C are needed to react with 0.500 mole of SO2? Answer: 1.25 mole of O2 But how??

OpenStudy (cyber405):

In reaction 2 moles of SO2 reacts with 5 moles of C... So for 0.5 mole of SO2 we should have the same ratio ... C/SO2 = 5/2 Is that clear?

OpenStudy (chmvijay):

2 mole of SO need 5 moles of carbon to react then 0.5 moles of So reacts needs =0.5*5 /2 =1.25 is that correct:)